A circle has a circumference of {12\pi}. It has an arc of length {16}{15}\pi

What is the central angle of the arc, in radians?

Answer by Alan3354(69443)

A circle has a circumference of {12\pi}. It has an arc of length {16}{15}\pi

What is the central angle of the arc, in radians?
=============
Is {12\pi} supposed to be 12/pi?
---
What does {16}{15}\pi mean?